OpenStudy (anonymous):

Use a linear approximation (or differentials) to estimate the given number: 1/4.002 I'm need help finding the steps

OpenStudy (anonymous):

0.25

OpenStudy (anonymous):

.... i think you can find that on calculator

OpenStudy (anonymous):

what if he didnt want to use the calc

OpenStudy (anonymous):

oh sorry. It's my first time using this. i need to find the linear approx of it

OpenStudy (anonymous):

I just dont know the steps

OpenStudy (anonymous):

thats fine bro.. u cool

OpenStudy (anonymous):

then that means he's lazy not saying you are and dayveed it's okay trindawg is just stupid

OpenStudy (anonymous):

oops did i say that?

jimthompson5910 (jim_thompson5910):

Hint: your function is f(x) = 1/x you need to find f ' (x) and evaluate f ' (4)

jimthompson5910 (jim_thompson5910):

the point (4, 1/4) lies on the graph of f(x) = 1/x you need to find the equation of the tangent line L(x) at this point

OpenStudy (anonymous):

So i got f(x) =1/16

jimthompson5910 (jim_thompson5910):

how are you getting that

OpenStudy (anonymous):

f'(x) **

jimthompson5910 (jim_thompson5910):

f ' (x) is not 1/16

OpenStudy (anonymous):

Don't i find f '(x) of 1/x?

OpenStudy (anonymous):

which is -1/x^2

OpenStudy (anonymous):

so if i plug in x=4 it would make it - 1/16

jimthompson5910 (jim_thompson5910):

yes f ' (x) = -1/x^2 so f ' (4) = -1/16

jimthompson5910 (jim_thompson5910):

the slope of the tangent line is m = -1/16

jimthompson5910 (jim_thompson5910):

the point this tangent line goes through is (4, 1/4)

OpenStudy (anonymous):

so would i do y - 1/4 = -1/16(x-4)?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

solve for y to determine L(x) then you can compute L(4.002) to approximate f(4.002)

OpenStudy (anonymous):

okay so y = -1/16x + 1/2 right? so i just plug in 4.002 to x and that's my answer?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Oooo wow nevermind i got it. My teacher taught me a different way so i was confused about where you were going. This way makes a lot more sense. thank you!

jimthompson5910 (jim_thompson5910):

sure thing